OpenStudy (anonymous):

Pls HELP! a person is walking towards a mountain. At one point, the angle of elevation from the ground to the top of the mountain is 37 degrees. After walking another 1000 feet, the angle of elevation from the ground to the top of the mountain is 40 degrees. how high is the mountain?

OpenStudy (anonymous):

I know ur suppose to use the law of sine..but idk how

OpenStudy (campbell_st):

draw a picture  label your known information and then find the angle supplementary to 40, 140 and include that. the 3rd angle in the triangle will be 180 - 37 = 140 now you need to find x use the law of sines \[\frac{x}{\sin(37)} = \frac{1000}{\sin(3)}\] when you have found x you have the hypotenuse of the right triangle and can find the height using the sin ratio. hope this helps, and makes some sense.

OpenStudy (anonymous):

@campbell_st Thank you so much!